Gucchi:

Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2, and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The following equation represents Sal's profits last month, 2x + 3y = 1,470 where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold. What is the Slope- Intercept form, y-intercept, and slope for this problem

Gucchi:

The y intercept is 490 and slop is -2x/3 right?

Gucchi:

what do i put for the slope intercept form

Gucchi:

@axie

XxXNessalulbaddieXxX:

hai hai

Gucchi:

@tetsxpreme

Gucchi:

@xxxnessalulbaddiexxx wrote:
hai hai
hi

XxXNessalulbaddieXxX:

ur question has been answered alr

XxXNessalulbaddieXxX:

https://questioncove.com/updates/5627f07ee4b07052cc599c9e asked and answered here

Gucchi:

yh but what about the slope intercept form

Gucchi:

3y = 1470 – 2x

Gucchi:

this?

XxXNessalulbaddieXxX:

@gucchi wrote:
yh but what about the slope intercept form
i belivve it rights out the problem for you unless it does not

Gucchi:

bro i just want to know the slope intercept form

XxXNessalulbaddieXxX:

@gucchi wrote:
bro i just want to know the slope intercept form
ight bro

carmelle:

@gucchi wrote:
3y = 1470 – 2x
no, that wouldn't be correct. your question is answered in the link that nessa gave but basically, slope intercept form is y = mx +b m = slope y = y-intercept You already know the slope and y-intercept so just plug in the values instead of the words :)

axie:

I always use some thing called desmos.com to find these out. pretty easy to use tool

Gucchi:

wait so y = 2x + 490

Gucchi:

@axie wrote:
I always use some thing called desmos.com to find these out. pretty easy to use tool
same its helps out alot

TETSXPREME:

@gucchi wrote:
wait so y = 2x + 490
no i dont think so

Gucchi:

so where does 3y go?

carmelle:

@gucchi wrote:
wait so y = 2x + 490
is your slope "2x" or `"-2/3x"`?

Gucchi:

OHH

Gucchi:

my b

Gucchi:

@carmelle wrote:
@gucchi wrote:
wait so y = 2x + 490
is your slope "2x" or `"-2/3x"`?
so y = -2/3x + b

Gucchi:

but does the x go after 2 or 3?

Gucchi:

or can it be either

TETSXPREME:

It goes after 2

carmelle:

@gucchi wrote:
@carmelle wrote:
@gucchi wrote:
wait so y = 2x + 490
is your slope "2x" or "-2/3x"?
so y = -2/3x + b
yea, you're almost there, remember, b = y-intercept

Gucchi:

yh my y inter is 490 right

carmelle:

@carmelle wrote:
@gucchi wrote:
3y = 1470 – 2x
no, that wouldn't be correct. your question is answered in the link that nessa gave but basically, slope intercept form is y = mx +b m = slope y = y-intercept You already know the slope and y-intercept so just plug in the values instead of the words :)
b = y-intercept* typo sorry

carmelle:

@gucchi wrote:
yh my y inter is 490 right
exactly, so your final answer would be what?

Gucchi:

y = -2x/3 + 490 but what abt 3y?

carmelle:

we solved for y so 3y is gone, do you understand?

Gucchi:

ohh true

Gucchi:

yh

carmelle:

great :)

Gucchi:

ty so much

carmelle:

yw